OpenStudy (anonymous):

Can someone please explain this to me I am stuck on this one question. Describe the differences between the graph of y = –3(x + 7)2 – 10 and the standard position graph of y = x2.

OpenStudy (anonymous):

@i0iz0gangster

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@Peasant1

OpenStudy (mathstudent55):

Compare \(y = a(x - b)^2 + c \) with \(y = x^2 \) \(a\) makes the parabola wider of narrower. \(b\) shifts the graph left or right \(c\) shifts the graph up or down.